# Deep neural networks for choice analysis: Enhancing behavioral regularity with gradient regularization

### v1

@@ Results
Firstly, the benchmark DNNs without gradient regularization outperforms the MNL models in predictive power but underperforms the MNL in behavioral regularity, especially with the Chicago data. The benchmark DNN improves the MNL’s log-likelihood by 5.2% of its absolute value with the Chicago data and 5.5% with the London data. The empirical results show that log-likelihood is more sensitive to predictive performance than accuracy and $F_{1}$ score due to its probabilistic nature. Meanwhile, the difference between accuracy and $F_{1}$ score reflects whether the dataset is balanced: the two metrics are similar for the London data but remain a gap for the car-dominated Chicago data, indicating the London data is more balanced. Moreover, the benchmark DNNs present significant behavioral irregularity, e.g., as suggested by the relatively lower strong regularity (88.8%) and weak regularity (92.2%) with the Chicago data. This is consistent with finding illustrated in Fig. 2a: the average market share of driving is non-monotonic w.r.t. driving cost, which is consistently presented in all local DNN models. This suggests that DNNs typically have high predictive performance but low behavioral regularity, aligning with the findings from many previous studies (Wang et al., 2020a, b; Wong and Farooq, 2021; Xia et al., 2023). By contrast, the benchmark TasteNet has higher behavioral regularity but lower predictive power, possibly because it has a simpler NN architecture and is linear in taste parameters.
@@
Secondly, sum-based gradient regularization can improve the behavioral regularity of DNNs and TasteNets without sacrificing their predictive power. The demand functions of the regularized DNNs become highly monotonic, as indicated by strong and weak regularities both approaching 1. Meanwhile, sum-based gradient regularization is able to enhance DNNs’ predictive power in both datasets. Table 2 demonstrates that sum-PGR, which directly regularizes the demand functions, is slightly more effective than sum-UGR and sum-LGR, which exploit demand monotonicity through the chain rule. This finding is further elaborated by Fig. 2b, where the regularized DNN has individual demand functions more consistent with the MNL: driving is less favored due to increased costs, while train and active mobility see higher demand. The regularized demand functions are more monotonic not only in the ensemble model, but also in training replications. We note that similar behavioral regularization effect of sum-based gradient regularization is also observed for TasteNets on the London dataset.
@@
On the other hand, norm-based gradient regularization fails to enhance behavioral regularity or predictive power. For example, norm-UGR with a small $\lambda$ can preserve accuracy and behavioral regularity for DNNs and TasteNets, but it would eventually lead to low predictive power and strong regularity as we increase $\lambda$. These results hold for all three norm-based approaches, potentially because they tend to flatten and smooth the demand curves. Therefore, the optimal DNNs and TasteNets in Table 2 have small $\lambda$’s and look similar to the corresponding benchmarks (see Fig. 2c for example). With strong norm-based gradient regularization, as shown in Fig. D.1e, individual demand curves become almost flat and could not reflect the decision mechanism: travelers might not respond to cost changes at certain points, but are highly unlikely to be insensitive to all cost changes. Since smaller $\lambda$’s lead to better performance, we might conclude that the datasets or models do not change abruptly due to cost perturbations. In brief, although regularizing the gradient norm is a common practice in computer science (Drucker and Le Cun, 1991; Jakubovitz and Giryes, 2018; Sokolić et al., 2017), it is not founded on prior beliefs in behavioral regularity and thus not effective for our purposes. For conciseness, we will not analyze norm-based gradient regularization in the next two scenarios.
@@
In this subsection, we focus on exemplifying the effectiveness of the proposed gradient regularizers in a typical choice modeling scenario where the number of available samples is limited due to resource limitations or privacy concerns. The same analysis is applied to the 1K-Random Chicago and London datasets. Table 3 illustrates the performance of DNNs, TasteNets, and benchmark models, while Fig. 3 elaborates on the individual demand functions. Further, the model performances in the training and validation sets are summarized in Tables C.3 and C.4, respectively.
We find that our gradient regularizers are even more effective than in the large sample scenario. Firstly, without GR, behaviorally regular become much worse in benchmark DNNs with small sample, e.g., the strong regularity dropped by 22.4 percentage points and the weak regularity dropped by 19.4 percentage points with the Chicago data. Similarly, shallower NNs like the benchmark TasteNets also have worse behavioral regularity in the small sample scenario. Meanwhile, the benchmark models fail to outperform MNL in predictive power, especially for the Chicago data. As shown in Fig. 3a, the benchmark DNN’s individual demand curves are non-monotonic and contradictory to the law of demand. Secondly, sum-based gradient regularization succeeds in enhancing all metrics of DNNs and most metrics of TasteNets, as shown in Table 3 and Fig. 3. In other words, the regularized DNNs and TasteNets outperform their corresponding benchmark models in both predictive power and behavioral regularity. With the Chicago data, for example, sum-PGR improves the benchmark DNN’s log-likelihood by 1.9% of its absolute value and its strong regularity by 32.1 percentage points. In addition, ensembles of regularized models have comparable overall performance to the MNL when using such a small sample size.
@@
mean (std.) No GR PGR UGR LGR No GR PGR UGR LGR MNL
Log-likelihood $-375.1$ $\mathbf{-367.9}$ $-374.2$ $-369.7$ $-389.3$ $-385.2$ $\underline{-383.1}$ $-386.6$ $-380.2$
(3.096) (4.181) (3.298) (3.304) (2.838) (3.351) (3.139) (3.447) (0)
Accuracy 0.705 0.703 0.676 0.696 0.700 0.679 0.686 0.677 0.718
(0.002) (0.007) (0.01) (0.011) (0.005) (0.012) (0.010) (0.012) (0)
$F_{1}$ score 0.648 0.645 0.559 0.619 0.619 0.563 0.580 0.559 0.665
(0.004) (0.018) (0.027) (0.030) (0.012) (0.028) (0.023) (0.028) (0)
Strong regularity 0.664 0.985 0.985 0.985 0.659 0.979 0.983 0.979 0.996
(0.173) (0.009) (0.011) (0.011) (0.129) (0.015) (0.013) (0.013) (0)
Weak regularity 0.728 0.999 0.997 0.999 0.685 0.997 0.995 0.992 1.000
(0.162) (0.001) (0.005) (0.002) (0.128) (0.004) (0.006) (0.009) (0)
Panel 2: London dataset, sum-XGR
Log-likelihood $-322.6$ $-328.8$ $\mathbf{-317.8}$ $-335.2$ $-345.1$ $\underline{-339.7}$ $-343.5$ $-340.1$ $-331.2$
(3.455) (8.146) (3.44) (11.304) (2.337) (2.023) (3.035) (1.974) (0)
Accuracy 0.737 0.720 0.740 0.717 0.725 0.734 0.724 0.733 0.746
(0.007) (0.013) (0.007) (0.021) (0.004) (0.006) (0.009) (0.006) (0)
$F_{1}$ score 0.732 0.703 0.734 0.694 0.711 0.720 0.705 0.718 0.740
(0.007) (0.020) (0.008) (0.038) (0.005) (0.006) (0.013) (0.005) (0)
Strong regularity 0.904 0.999 0.992 0.997 0.939 0.964 0.999 0.965 0.998
(0.069) (0.002) (0.014) (0.008) (0.023) (0.013) (0.002) (0.01) (0)
Weak regularity 0.913 1.000 0.993 0.998 0.943 0.984 0.999 0.981 1.000
(0.067) (0.001) (0.014) (0.006) (0.024) (0.012) (0.002) (0.013) (0)
@@
(c) TasteNet
(d) TasteNet, sum-PGR
(e) MNL
Figure 3: Individual demands as functions of driving cost (1K-Random, Chicago).
@@
The large and small sample scenarios above assist in examining only in-domain generalization for the random split of training and test data. Although random split is the most common practice, we are also interested in the out-of-domain generalizability of DNNs. Out-of-domain generalization is highly relevant to transportation engineering, system design, and urban planning, such as the cross-city policy transfer: when assessing whether a city should build a subway, transportation planners sometimes cite the ridership of subway systems in other cities. Our data split scheme in the 10K-Sorted sub-datasets can be interpreted as using the patterns of individual choice behavior in one city (as in the training set) to extrapolate those in another city (as in the test set). Here we emulate such policy setting by testing the DNNs’ out-of-domain generalizability using predictive power and behavioral regularity metrics.
The results suggest that gradient regularization could drastically improve the out-of-domain generalizability of DNNs, even more effectively than improving their in-domain generalizability. Table 4 summarizes the performance of DNNs, TasteNets, and benchmark models in the test sets, while their training and validation performance are summarized in Tables C.5 and C.6. Under this setting, DNNs and TasteNets exhibit great flexibility with higher log-likelihood than the MNL model. However, behavioral regularity could be relatively low for these NN models, as reflected in their regularity metrics. Fig. 4 visualizes the individual demand functions as functions of transit cost with the London data, in which the benchmark DNN performs unreasonably in the test set shown to the right of the data split threshold (dashed gray line). Secondly, sum-based gradient regularization has the potential to simultaneously improve the predictive power and behavioral regularity of benchmark DNNs and TasteNets. For an extreme case, the benchmark DNN with the London data, sum-UGR dramatically raises its strong regularity from 0.185 to 0.979. However, sum-based gradient regularization is unable to further improve the benchmark TasteNet with the Chicago data, whose base behavioral regularity is high.
@@
mean (std.) No GR PGR UGR LGR No GR PGR UGR LGR MNL
Log-likelihood $-1356.1$ $-1243.5$ $\mathbf{-1167.9}$ $-1232.4$ $\underline{-1485.9}$ $-1873.7$ $-1901.4$ $-2003.9$ $-2025.7$
(134.1) (55.1) (27.6) (54.9) (47.9) (45.5) (36.6) (50.2) (0)
Accuracy 0.783 0.788 0.789 0.788 0.750 0.726 0.725 0.705 0.722
(0.011) (0.002) (0.003) (0.002) (0.014) (0.008) (0.006) (0.008) (0)
$F_{1}$ score 0.722 0.726 0.727 0.724 0.707 0.721 0.720 0.708 0.721
(0.010) (0.006) (0.003) (0.005) (0.006) (0.005) (0.004) (0.006) (0)
Strong regularity 0.317 0.857 0.923 0.865 1.000 0.980 0.978 0.981 0.984
(0.240) (0.099) (0.087) (0.071) (0.000) (0.004) (0.004) (0.004) (0)
Weak regularity 0.487 0.974 0.983 0.977 1.000 1.000 1.000 1.000 1.000
(0.230) (0.037) (0.040) (0.023) (0.000) (0.000) (0.000) (0.001) (0)
Panel 2: London dataset, sum-XGR (transit cost)
Log-likelihood $-1137.9$ $-1110.4$ $\mathbf{-1108.4}$ $-1112.1$ $-1221.1$ $\underline{-1170.2}$ $-1171.1$ $-1175.0$ $-1301.6$
(34.3) (23.9) (21.1) (26.7) (33.1) (27.0) (28.5) (30.3) (0)
Accuracy 0.777 0.784 0.794 0.782 0.776 0.786 0.785 0.783 0.780
(0.010) (0.007) (0.007) (0.006) (0.009) (0.008) (0.008) (0.007) (0)
$F_{1}$ score 0.768 0.776 0.778 0.776 0.765 0.772 0.772 0.772 0.770
(0.011) (0.006) (0.008) (0.006) (0.008) (0.008) (0.008) (0.007) (0)
Strong regularity 0.185 0.968 0.979 0.907 0.313 0.878 0.872 0.720 0.980
(0.075) (0.032) (0.029) (0.062) (0.007) (0.095) (0.095) (0.083) (0)
Weak regularity 0.207 0.977 0.984 0.925 0.343 0.993 0.988 0.900 1.000
(0.080) (0.026) (0.025) (0.055) (0.008) (0.010) (0.012) (0.045) (0)
@@
(c) TasteNet
(d) TasteNet, sum-PGR
(e) MNL
Figure 4: Individual demands as functions of transit cost (10K-Sorted, London).
@@
In a large sample scenario, higher behavioral regularity often has a trade-off with lower predictive power. The effects of regularization strength $\lambda$ on the five metrics are illustrated in Fig. 5 for both 10K-Random Chicago and London datasets, where the horizontal axis uses logarithmic scale $\lg(\lambda)$. As shown in Fig. 5, DNNs’ model fit declines with increasing $\lambda$, especially in terms of log-likelihood, which is more sensitive to $\lambda$ than accuracy and $F_{1}$ score. The decline in predictive power is particularly noticeable after a critical point, such as $\lambda=10$ for DNNs with sum-PGR. In other words, although regularization would generally reduce predictive power, there exists a range of $\lambda$ that almost preserves predictive power while enhancing behavioral regularity, such as $\lambda\leq 0.1$ for DNNs with sum-PGR, consistent with our findings in Section 5.1.1.
(a) DNN, sum-PGR (Chicago)
(b) DNN, norm-PGR (Chicago)
(c) DNN, sum-PGR (London)
(d) TasteNet, sum-PGR (Chicago)
(e) TasteNet, norm-PGR (Chicago)
(f) TasteNet, sum-PGR (London)
@@
Fig. 5 also presents two important differences between log-likelihood and accuracy or $F_{1}$ score, as well as between sum- and norm-based gradient regularization. Firstly, log-likelihood is much more sensitive than accuracy or $F_{1}$ score in measuring predictive power, which is theoretically valid and empirically expected. Moreover, presenting accuracy and $F_{1}$ score together might be a good idea for imbalanced data. Therefore, we recommend that future studies use different metrics when comparing the performance of DCMs. Secondly, our behavioral regularity metrics can demonstrate the flattening effects of norm-based gradient regularization on individual demand curves, as indicated by weak regularity approaching 1 and strong regularity approaching 0 for very large $\lambda$’s (see Fig. 5b). In brief, strong regularity might be more appropriate than the weak one, at least for describing the global declining trend of demand curves, although weak behavioral regularity metric could still be important for describing local insensitivity to cost changes.
@@
Interestingly, under a setting with relatively small sample, stronger sum-based gradient regularization can simultaneously improve predictive power and behavioral regularity, thereby advancing the Pareto frontier of different performance metrics. Fig. 6 illustrates the effects of $\lambda$ for the 1K-Random sub-datasets, where we can find a range for each gradient regularizer such that predictive power and behavioral regularity increase together. For example, with the Chicago data, when $\lambda$ increases from $10^{-4}$ to 0.1, sum-PGR improves the DNN’s log-likelihood by 1.9% of its absolute value and strong regularity by 29.4 percentage points. This phenomenon suggests the presence of Pareto efficiency in NN models, despite the substitution effects observed in Fig. 5.
(a) DNN, sum-PGR (Chicago)
(b) DNN, sum-PGR (London)
(c) TasteNet, sum-PGR (Chicago)
@@
On the other hand, when $\lambda$ exceeds a certain critical point, we still observe substitution effects between predictive power and behavioral regularity for both DNNs and TasteNets. For example, with the Chicago data, when $\lambda$ increases from 0.1 to 100, the sum-UGR reduces the DNN’s log-likelihood by 4.8% of its absolute value, but still improves the DNN’s strong regularity by 2.0 percentage points. This also demonstrates the flexibility of soft constraints by identifying the optimal $\lambda$, which reflects the alignment between the data and our behavioral assumptions.
@@
To explore the out-of-domain generalizability of DNNs with sum-based gradient regularization, we visualize the effects of $\lambda$ for the 10K-Sorted sub-datasets in Fig. 7. Interestingly, the trend of each metric combines the characteristics of the first two scenarios: we observe both complementary and substitution effects for DNNs, whereas substitution effects are more significant for TasteNets. In addition, the log-likelihood patterns for DNNs with the Chicago and London data are slightly different (see Fig. 7a and b), suggesting the data-dependency nature of NN training. On the other hand, consistent with findings from previous sections, complementary effects and Pareto efficiency are observed for the sum-based gradient regularizers. These results imply the usefulness of gradient regularization in improving behavioral regularity, as well as predictive power simultaneously for the neural network choice models.
(a) DNN, sum-UGR (Chicago)
(b) DNN, sum-UGR (London)
(c) TasteNet, sum-PGR (London)
@@
As a summary, our findings demonstrate the conditions under which our sum-based gradient regularizers and behavioral regularity metrics are most effective. When the sample size is large, DNN models tend to have a high behavioral regularity. In this case, even a benchmark DNN architecture could be used for behavioral prediction and mobility policy analysis, so our behavioral regularity metrics and regularization methods might not be necessary. However, since it is always costly to collect travel survey data, the sample size in travel demand modeling is typically small (e.g., less than 10,000 samples). Under this context, sum-based gradient regularization should be broadly applied since it enables DNNs to generate reliable prediction and intuitive behavioral interpretation. Gradient regularization is also effective when the models are applied to future forecasting or cross-city transfer of mobility policies, which is emulated by the out-of-domain generalization scenario in our experiments.

### v2

@@ Results
First, the benchmark DNNs without gradient regularization outperform the MNL models in predictive power but underperform them in behavioral regularity, especially for the CMAP data. The benchmark DNN improves the MNL’s log-likelihood by 5.2% of its absolute value using the CMAP data and 5.5% using the LTDS data. The empirical results show that log-likelihood is more sensitive to predictive performance than accuracy and $F_{1}$ score due to its probabilistic nature. Meanwhile, the difference between accuracy and $F_{1}$ score reflects whether the dataset is balanced: the two metrics are similar for the LTDS data but remain a gap for the car-dominated CMAP data, indicating that the former is more balanced in travel modes. Moreover, the benchmark DNNs are behaviorally irregular, as suggested by the relatively lower strong regularity (88.8%) and weak regularity (92.2%) using the CMAP data. This is consistent with finding illustrated in Fig. 2a: the average market share of driving is non-monotonic w.r.t. driving cost, which is consistently presented in all local DNN models. This suggests that DNNs typically have high predictive performance but low behavioral regularity, aligning with the findings from many previous studies (Wang et al., 2020a, b; Wong and Farooq, 2021; Xia et al., 2023). By contrast, the benchmark TasteNets have high behavioral regularity but relatively low predictive power, possibly due to their simpler NN architecture and linear-in-parameters specifications.
Second, the DNNs with soft constraints outperform the TasteNets with hard constraints in terms of log-likelihood. By design, both ReLU and exponential constraints guarantee 100% weak regularity, but only the latter can simultaneously guarantee high strong regularity. Despite being behaviorally regular, the TasteNets with either soft or hard constraints have lower log-likelihood than the DNNs with soft constraints. This example demonstrates the flexibility and approximation power of nonlinear DNN specifications compared with the linear-in-parameters TasteNet specifications. It might also suggest the existence of irregular behaviors for certain individuals or price ranges, highlighting the flexibility of soft constraints in capturing various behavioral mechanisms (Bagwell and Riordan, 1991; Stiving, 2000). As shown in Tables 2 and 2, the exponentially constrained TasteNets resemble the MNL models in performance metrics and individual demands, reflecting the intrinsic connection between TasteNet and MNL models.
@@
Third, sum-based gradient regularization can improve the behavioral regularity of DNNs and TasteNets without sacrificing their predictive power. The demand functions of the regularized DNNs become highly monotonic, as indicated by strong and weak regularities both approaching 100%. Meanwhile, sum-based gradient regularization is able to enhance DNNs’ predictive power in both datasets. Table 2 demonstrates that sum-PGR, which directly regularizes the demand functions, is slightly more effective than sum-UGR and sum-LGR, which exploit demand monotonicity through the chain rule. This finding is further elaborated by Fig. 2b, where the regularized DNN has individual demand functions more consistent with the MNL: driving is less favored due to increased costs, while train and active mobility see higher demand. The regularized demand functions are more monotonic not only in the ensemble model, but also in training replications. We note that similar behavioral regularization effects of sum-based gradient regularizers are also observed for TasteNets on the LTDS data.
@@
Last but not least, norm-based gradient regularization fails to enhance behavioral regularity or predictive power. For example, norm-UGR with a small $\lambda$ can preserve accuracy and behavioral regularity for DNNs and TasteNets, but it would eventually lead to low predictive power and strong regularity as we increase $\lambda$. These results hold for all three norm-based approaches, potentially because they tend to flatten and smooth the demand curves. Therefore, the optimal DNNs and TasteNets in Table 2 have small $\lambda$’s and look similar to the corresponding benchmarks (see Fig. 2c for example). With strong norm-based gradient regularization, as shown in Fig. 10e, individual demand curves become almost flat and could not reflect the decision mechanism: travelers might not respond to cost changes at certain points, but are highly unlikely to be insensitive to all cost changes. Since smaller $\lambda$’s lead to better performance, we might conclude that the datasets or models do not change abruptly due to cost perturbations. In brief, although regularizing the gradient norm is a common practice in computer science (Drucker and Le Cun, 1991; Jakubovitz and Giryes, 2018; Sokolić et al., 2017), it is not founded on prior beliefs in behavioral regularity and thus not effective for our purposes. For conciseness, we will not analyze norm-based gradient regularization in the next two scenarios.
@@
In this subsection, we focus on exemplifying the effectiveness of the proposed gradient regularizers in a typical choice modeling scenario where the number of available samples is limited due to resource restrictions or privacy concerns. The same analysis is applied to the 1K-Random CMAP and LTDS datasets. Table 3 illustrates the performance of DNNs, TasteNets, and benchmark models, while Fig. 3 elaborates on the individual demand functions. Furthermore, Section C.2 summarizes the training and validation performance of the models.
@@
Mean (SD) No GR PGR UGR LGR No GR PGR UGR LGR ReLU Exp MNL
Panel 1: CMAP data, sum-XGR
Log-likelihood $-375.1$ $\mathbf{-367.9}$ $-374.2$ $\underline{-369.7}$ $-389.3$ $-385.2$ $-383.1$ $-386.6$ $-386.3$ $-378.0$ $-380.2$
(3.096) (4.181) (3.298) (3.304) (2.838) (3.351) (3.139) (3.447) (2.506) (4.123) (0)
Accuracy 0.705 0.703 0.676 0.696 0.700 0.679 0.686 0.677 0.696 0.707 0.718
(0.002) (0.007) (0.010) (0.011) (0.005) (0.012) (0.010) (0.012) (0.005) (0.004) (0)
$F_{1}$ score 0.648 0.645 0.559 0.619 0.619 0.563 0.580 0.559 0.610 0.637 0.665
(0.004) (0.018) (0.027) (0.030) (0.012) (0.028) (0.023) (0.028) (0.011) (0.010) (0)
Strong regularity 0.664 0.985 0.985 0.985 0.659 0.979 0.983 0.979 0.461 0.999 0.996
(0.173) (0.009) (0.011) (0.011) (0.129) (0.015) (0.013) (0.013) (0.195) (0.001) (0)
Weak regularity 0.728 0.999 0.997 0.999 0.685 0.997 0.995 0.992 1.000 1.000 1.000
(0.162) (0.001) (0.005) (0.002) (0.128) (0.004) (0.006) (0.009) (0.000) (0.000) (0)
Panel 2: LTDS data, sum-XGR
Log-likelihood $\underline{-322.6}$ $-328.8$ $\mathbf{-317.8}$ $-335.2$ $-345.1$ $-339.7$ $-343.5$ $-340.1$ $-341.0$ $-335.5$ $-331.2$
(3.455) (8.146) (3.440) (11.30) (2.337) (2.023) (3.035) (1.974) (2.572) (2.951) (0)
Accuracy 0.737 0.720 0.740 0.717 0.725 0.734 0.724 0.733 0.729 0.733 0.746
(0.007) (0.013) (0.007) (0.021) (0.004) (0.006) (0.009) (0.006) (0.006) (0.008) (0)
$F_{1}$ score 0.732 0.703 0.734 0.694 0.711 0.720 0.705 0.718 0.715 0.725 0.740
(0.007) (0.020) (0.008) (0.038) (0.005) (0.006) (0.013) (0.005) (0.007) (0.009) (0)
Strong regularity 0.904 0.999 0.992 0.997 0.939 0.964 0.999 0.965 0.924 0.999 0.998
(0.069) (0.002) (0.014) (0.008) (0.023) (0.013) (0.002) (0.01) (0.046) (0.002) (0)
Weak regularity 0.913 1.000 0.993 0.998 0.943 0.984 0.999 0.981 1.000 1.000 1.000
(0.067) (0.001) (0.014) (0.006) (0.024) (0.012) (0.002) (0.013) (0.000) (0.000) (0)
We find that our gradient regularizers are even more effective than in the large sample scenario. First, without gradient regularization, the benchmark DNNs’ behavioral regularity becomes much worse with a small sample, e.g., the strong regularity drops by 22.4 percentage points and the weak regularity drops by 19.4 percentage points using the CMAP data. Similarly, shallower NNs like the benchmark TasteNets also have worse behavioral regularity in the small sample scenario. Meanwhile, the benchmark models fail to outperform MNL in predictive power, especially for the CMAP data. As shown in Fig. 3a, the benchmark DNN’s individual demand curves are non-monotonic and contradictory to the law of demand. Second, sum-based gradient regularization succeeds in enhancing all metrics of DNNs and most metrics of TasteNets, as shown in Table 3 and Fig. 3. In other words, the regularized DNNs and TasteNets outperform their corresponding benchmark models in both predictive power and behavioral regularity. In the CMAP case, sum-PGR improves the benchmark DNN’s log-likelihood by 1.9% of its absolute value and its strong regularity by 32.1 percentage points. Moreover, from Tables 3 and 3 we can draw similar findings for the hard constraints, e.g., the exponentially constrained TasteNets resemble the MNL models in performance metrics and individual demands. In addition, ensembles of regularized models have comparable overall performance to the MNL when using such a small sample size.
@@
(c) MNL
(d) TasteNet
(e) TasteNet, sum-PGR
(f) TasteNet, Exp
Figure 3: Individual demands as functions of driving costs (1K-Random, CMAP).
@@
The large and small sample scenarios above assist in examining only in-domain generalization for the random split of training and test data. Although random split is the most common practice, we are also interested in the out-of-domain generalizability of DNNs. Out-of-domain generalization is highly relevant to transportation engineering, system design, and urban planning, such as the cross-city policy transfer: when assessing whether a city should build a subway, transportation planners sometimes cite the ridership of subway systems in other cities. Our data split scheme in the 10K-Sorted datasets can be interpreted as using the patterns of individual choice behavior in one city (as in the training set) to extrapolate those in another city (as in the test set). Here we emulate such policy setting by testing the DNNs’ out-of-domain generalizability using predictive power and behavioral regularity metrics.
@@
Mean (SD) No GR PGR UGR LGR No GR PGR UGR LGR ReLU Exp MNL
Panel 1: CMAP data, sum-XGR
Log-likelihood $-1356.1$ $-1243.5$ $\mathbf{-1167.9}$ $\underline{-1232.4}$ $-1485.9$ $-1873.7$ $-1901.4$ $-2003.9$ $-2315.4$ $-1995.0$ $-2025.7$
(134.1) (55.15) (27.64) (54.87) (47.92) (45.52) (36.62) (50.23) (395.4) (51.83) (0)
Accuracy 0.783 0.788 0.789 0.788 0.750 0.726 0.725 0.705 0.578 0.723 0.722
(0.011) (0.002) (0.003) (0.002) (0.014) (0.008) (0.006) (0.008) (0.134) (0.007) (0)
$F_{1}$ score 0.722 0.726 0.727 0.724 0.707 0.721 0.720 0.708 0.603 0.719 0.721
(0.010) (0.006) (0.003) (0.005) (0.006) (0.005) (0.004) (0.006) (0.112) (0.005) (0)
Strong regularity 0.317 0.857 0.923 0.865 1.000 0.980 0.978 0.981 0.933 0.969 0.984
(0.240) (0.099) (0.087) (0.071) (0.000) (0.004) (0.004) (0.004) (0.169) (0.005) (0)
Weak regularity 0.487 0.974 0.983 0.977 1.000 1.000 1.000 1.000 1.000 1.000 1.000
(0.230) (0.037) (0.040) (0.023) (0.000) (0.000) (0.000) (0.001) (0.000) (0.000) (0)
Panel 2: LTDS data, sum-XGR (public transit cost)
Log-likelihood $-1137.9$ $\underline{-1110.4}$ $\mathbf{-1108.4}$ $-1112.1$ $-1221.1$ $-1170.2$ $-1171.1$ $-1175.0$ $-1182.3$ $-1150.1$ $-1301.6$
(34.33) (23.86) (21.15) (26.67) (33.10) (27.01) (28.54) (30.25) (54.11) (29.62) (0)
Accuracy 0.777 0.784 0.794 0.782 0.776 0.786 0.785 0.783 0.782 0.781 0.780
(0.010) (0.007) (0.007) (0.006) (0.009) (0.008) (0.008) (0.007) (0.017) (0.008) (0)
$F_{1}$ score 0.768 0.776 0.778 0.776 0.765 0.772 0.772 0.772 0.769 0.768 0.770
(0.011) (0.006) (0.008) (0.006) (0.008) (0.008) (0.008) (0.007) (0.018) (0.008) (0)
Strong regularity 0.185 0.968 0.979 0.907 0.313 0.878 0.872 0.720 0.248 0.982 0.980
(0.075) (0.032) (0.029) (0.062) (0.007) (0.095) (0.095) (0.083) (0.084) (0.005) (0)
Weak regularity 0.207 0.977 0.984 0.925 0.343 0.993 0.988 0.900 1.000 1.000 1.000
(0.080) (0.026) (0.025) (0.055) (0.008) (0.010) (0.012) (0.045) (0.000) (0.000) (0)
@@
(c) MNL
(d) TasteNet
(e) TasteNet, sum-PGR
(f) TasteNet, Exp
Figure 4: Individual demands as functions of public transit costs (10K-Sorted, LTDS).
The results suggest that gradient regularization could drastically improve the out-of-domain generalizability of DNNs, even more effectively than improving their in-domain generalizability. Table 4 summarizes the performance of DNNs, TasteNets, and benchmark models in the test sets, while their training and validation performance are summarized in Section C.3. Under this setting, DNNs and TasteNets exhibit great flexibility with higher log-likelihood than the MNL model. However, behavioral regularity could be relatively low for these NN models, as reflected in their regularity metrics. Fig. 4 visualizes individual demands as functions of public transit costs using the LTDS data, in which the benchmark DNN performs unreasonably in the test set, as shown to the right of the data split threshold (dashed gray line). Second, sum-based gradient regularization has the potential to simultaneously improve the predictive power and behavioral regularity of benchmark DNNs and TasteNets. For an extreme case, sum-UGR dramatically raises the strong regularity of the benchmark DNN from 0.185 to 0.979 using the LTDS data. Table 4 also reveals the case-dependent effects of hard constraints: they can significantly improve both behavioral regularity and predictive power of a poorly behaved benchmark TasteNet (the LTDS case), but are ineffective for a benchmark that is already highly regular (the CMAP case).
@@
In a large sample scenario, higher behavioral regularity often has a trade-off with lower predictive power. The effects of regularization strength $\lambda$ on the five metrics are illustrated in Fig. 5 for both 10K-Random CMAP and LTDS datasets, where the horizontal axis uses logarithmic scale $\lg(\lambda)$. As shown in Fig. 5, DNNs’ model fit declines with increasing $\lambda$, especially in terms of log-likelihood, which is more sensitive to $\lambda$ than accuracy and $F_{1}$ score. The decline in predictive power is particularly noticeable after a critical point, such as $\lambda=10$ for DNNs with sum-PGR. In other words, although regularization would generally reduce predictive power, there exists a range of $\lambda$ that almost preserves predictive power while enhancing behavioral regularity, such as $\lambda\leq 0.1$ for DNNs with sum-PGR, consistent with our findings in Section 5.1.1.
(a) DNN, sum-PGR (CMAP)
(b) DNN, norm-PGR (CMAP)
(c) DNN, sum-PGR (LTDS)
(d) TasteNet, sum-PGR (CMAP)
(e) TasteNet, norm-PGR (CMAP)
(f) TasteNet, sum-PGR (LTDS)
@@
Fig. 5 also presents two important differences between log-likelihood and accuracy or $F_{1}$ score, as well as between sum- and norm-based gradient regularization. First, log-likelihood is much more sensitive than accuracy or $F_{1}$ score in measuring predictive power, which is theoretically valid and empirically expected. Moreover, presenting accuracy and $F_{1}$ score together might be a good idea for imbalanced data. Therefore, we recommend that future studies use different metrics when comparing the performance of DCMs. Second, our behavioral regularity metrics can demonstrate the flattening effects of norm-based gradient regularization on individual demand curves, as indicated by weak regularity approaching 100% and strong regularity approaching 0 for very large $\lambda$’s (see Fig. 5b). In brief, strong regularity might be more appropriate than the weak one, at least for describing the global declining trend of demand curves, although weak behavioral regularity metric could still be important for describing local insensitivity to cost changes.
@@
Interestingly, under a setting with relatively small sample, stronger sum-based gradient regularization can simultaneously improve predictive power and behavioral regularity, thereby advancing the Pareto frontier of different performance metrics. Fig. 6 illustrates the effects of $\lambda$ on the 1K-Random datasets, where we can find a range for each gradient regularizer such that predictive power and behavioral regularity increase together. In the CMAP case, when $\lambda$ increases from $10^{-4}$ to 0.1, sum-PGR improves the DNN’s log-likelihood by 1.9% of its absolute value and strong regularity by 29.4 percentage points. This phenomenon suggests the presence of Pareto efficiency in NN models, despite the substitution effects observed in Fig. 5.
(a) DNN, sum-PGR (CMAP)
(b) DNN, sum-PGR (LTDS)
(c) TasteNet, sum-PGR (CMAP)
@@
On the other hand, when $\lambda$ exceeds a certain critical point, we still observe substitution effects between predictive power and behavioral regularity for both DNNs and TasteNets. In the CMAP case, when $\lambda$ increases from 0.1 to 100, sum-UGR reduces the DNN’s log-likelihood by 4.8% of its absolute value, but still improves the DNN’s strong regularity by 2.0 percentage points. This also demonstrates the flexibility of soft constraints by identifying the optimal $\lambda$, which reflects the alignment between the data and our behavioral assumptions.
@@
To explore the out-of-domain generalizability of DNNs with sum-based gradient regularization, we visualize the effects of $\lambda$ for the 10K-Sorted datasets in Fig. 7. Interestingly, the trend of each metric combines the characteristics of the first two scenarios: we observe both complementary and substitution effects for DNNs, whereas substitution effects are more significant for TasteNets. In addition, the log-likelihood patterns for DNNs using the CMAP and LTDS data are slightly different (see Fig. 7a and b), suggesting the data-dependency nature of NN training. On the other hand, consistent with findings from previous sections, complementary effects and Pareto efficiency are observed for the sum-based gradient regularizers. These results imply the effectiveness of gradient regularization in improving behavioral regularity and predictive power simultaneously for the NN-based choice models.
(a) DNN, sum-UGR (CMAP)
(b) DNN, sum-UGR (LTDS)
(c) TasteNet, sum-PGR (LTDS)
@@
As a summary, our findings demonstrate the conditions under which our sum-based gradient regularizers and behavioral regularity metrics are most effective. When the sample size is large, DNN models tend to have a high behavioral regularity. In this case, even a benchmark DNN architecture could be used for behavioral prediction and mobility policy analysis, and our behavioral regularity metrics and regularization methods might not be necessary. However, since it is always costly to collect travel survey data, the sample size in travel demand modeling is typically small (e.g., less than 10,000 samples). Under this context, sum-based gradient regularization should be broadly applied since it enables DNNs to generate reliable prediction and intuitive behavioral interpretation. Gradient regularization is also effective when the models are applied to future forecasting or cross-city transfer of mobility policies, which is emulated by the out-of-domain generalization scenario in our experiments. More generally, soft and hard constraints have their own advantages in improving behavioral regularity and predictive power, but neither is necessarily beneficial if the benchmark model already exhibits high behavioral regularity.